OpenStudy (anonymous):

What is the slope between the two points (6, 20) and (2, 4)? 1 over 4 negative 1 over 4 4 -4 @jim_thompson5910

OpenStudy (karatechopper):

y1-y2/x1-x2=m

OpenStudy (anonymous):

http://sketchtoy.com/32894540 Is this right? @karatechopper

OpenStudy (bakonloverk):

20-4= 6-2= then put the first answer over the second answer and simplify

OpenStudy (bakonloverk):

better? @karatechopper

OpenStudy (bakonloverk):

@unheard your fraction is the wrong way

OpenStudy (anonymous):

4/16 ?

OpenStudy (bakonloverk):

yes and then simplify

OpenStudy (anonymous):

1/4?

OpenStudy (bakonloverk):

yep